# Secure numerical simulations using fully homomorphic encryption

### v1

@@ Abstract
Data privacy is a significant concern in many environments today. This is particularly true if sensitive information, e.g., engineering, medical, or financial data, is to be processed on potentially insecure systems, as it is often the case in cloud computing. Fully homomorphic encryption (FHE) offers a potential solution to this problem, as it allows for secure computations on encrypted data. In this paper, we investigate the viability of using FHE for privacy-preserving numerical simulations of partial differential equations. We first give an overview of the CKKS scheme, a popular FHE method for computations with real numbers. This is followed by an introduction of our Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a Julia wrapper for the C++ library OpenFHE and offer a user-friendly interface for secure arithmetic operations. We then present a performance analysis of the CKKS scheme within OpenFHE, focusing on the error and efficiency of different FHE operations. Finally, we demonstrate the application of FHE to secure numerical simulations by implementing two finite difference schemes for the linear advection equation using the SecureArithmetic.jl package. Our results show that FHE can be used to perform cryptographically secure numerical simulations, but that the error and efficiency of FHE operations must be carefully considered when designing applications.
@@ 1 Introduction
With the ongoing rapid increase of cloud computing [1] in many areas of industry, commercial services, or science, data security and data privacy are growing concerns for many practitioners and researchers. As more and more data is being processed remotely, the impact of a potential data breach becomes higher, rendering it even more attractive for a malicious actor to try to extract confidential information. While some highly sensitive data may also be handled locally, there are also use cases where this is not possible, e.g., when the algorithms used for processing are proprietary themselves. A well-established practice is thus the use of strong encryption technologies, which protects such data from unauthorized access. However, classical encryption algorithms can only secure data at rest or while in transit. For the actual processing, the data needs to be decrypted first, opening the door for potential attacks.
@@
In this paper, we aim to address these issues by first providing a brief overview of the CKKS scheme from the perspective of a computational scientist. We will then present two packages we developed in the Julia programming language, OpenFHE.jl and SecureArithmetic.jl, that allow one to easily experiment with novel FHE applications. Finally, we will analyze the performance and accuracy of basic FHE operations, and then demonstrate the first prototypes of FHE-secured numerical simulations by implementing two well-known numerical methods for solving partial differential equations. We will find that while secure numerical simulations through FHE are indeed possible, the computational overhead is currently prohibitive for many practical scenarios, and the numerical errors introduced by the encryption techniques must be carefully considered when designing applications. All code used in this paper, as well as the data generated during the experiments, is available in our reproducibility repository [17].
The remainder of the paper is structured as follows: In Sec. 2, we give an overview of the CKKS scheme for FHE with real numbers. Moreover, we introduce the open-source C++ library OpenFHE and the Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a convenient interface for secure arithmetic operations in Julia for experimentation and rapid prototyping. We discuss the implementation of the CKKS scheme in OpenFHE in Sec. 3, and analyze the performance and accuracy of basic FHE operations. In Sec. 4, we implement several finite difference schemes to solve the linear advection equation using FHE, and analyze the performance and accuracy of the cryptographically secured numerical solvers in Sec. 5. Finally, we summarize our findings in Sec. 6 and provide an outlook on future work.
@@ 6 Conclusions and outlook
Next, we implemented the first-order upwind and second-order Lax-Wendroff schemes for the linear advection equations for secure numerical simulations using the SecureArithmetic.jl package. For this purpose, we introduced a $\operatorname{circshift}$ operation that is compatible with the CKKS scheme and that works for both 1D and 2D data representations. We then proceeded to conduct a convergence test that verified the correctness of our implementations and showed that the CKKS encryption does not break the convergence properties of the finite difference schemes.
@@
Finally, we conducted homomorphically encrypted numerical simulations by building upon the tools and methods developed in the previous sections. Our results illustrate that it is feasible to implement secure numerical simulations with the CKKS scheme, with results that match the unencrypted simulations well. However, the encrypted operations introduce a large computational overhead and are a significant source of error, to the point that we were able to achieve less than single-precision floating point accuracy at the cost of a six-orders-of-magnitude increase in runtime. Nearly all of these undesirable effects can be attributed to the bootstrapping procedure, which is the most inaccurate and slowest operation in the CKKS scheme. However, bootstrapping is also necessary to refresh the multiplicative depth of the ciphertexts, making it indispensable for long-running simulations. Moreover, iterative bootstrapping can further improve the accuracy of an FHE algorithm, though at a non-negligible runtime penalty that needs to be taken into account. We also showed that the CKKS scheme implementation in OpenFHE is parallelizable, and that using multiple threads can lead to a significant reduction in the solution time.
@@
To summarize, our work demonstrates that it is possible to perform secure numerical simulations using fully homomorphic encryption. At the same time, the large computational overhead and precision loss in the bootstrapping operation are significant obstacles that need to be taken into account. Therefore, currently FHE-secured numerical simulations only make practical sense in specialized applications where the security of the data is of utmost importance. Thus, more detailed investigations are necessary to optimize the use of the CKKS scheme for numerical simulations. By providing a user-friendly interface to the OpenFHE library in Julia, and by implementing some important building blocks such as the $\operatorname{circshift}$ algorithms, we hope to lay the foundation for further research in this direction.
@@
In the future, we plan to extend our work in several areas. We would like to increase the complexity of the simulations by looking at systems of equations and at three-dimensional problems. This will also necessitate an extension of the $\operatorname{circshift}$ operation to support multiple ciphertexts and tensors. Furthermore, we will investigate the discretization of nonlinear equations, which will require us to handle additional arithmetic operations such as division and exponentiation. We also plan to evaluate the potential of other homomorphic encryption schemes, such as using the BFV scheme for integers together with fixed-point arithmetic. Moreover, we would like to better understand the observed behavior that the accuracy of secure numerical simulations with bootstrapping can improve over time. Finally, we plan to continue developing the SecureArithmetic.jl package as a tool for experimentation and rapid prototyping with encrypted algorithms, to open up FHE to a wider audience in the computational science community.
@@ Acknowledgments
We gratefully acknowledge the computing time provided by Numerical Simulation research group of Gregor Gassner, University of Cologne. Michael Schlottke-Lakemper further acknowledges funding through the DFG research unit FOR 5409 "Structure-Preserving Numerical Methods for Bulk- and Interface Coupling of Heterogeneous Models (SNuBIC)" (project number 463312734), as well as through a DFG individual grant (project number 528753982).

### v2

@@ Abstract
Data privacy is a significant concern when using numerical simulations for sensitive information such as medical, financial, or engineering data. This issue becomes especially relevant in untrusted environments like public cloud infrastructures. Fully homomorphic encryption (FHE) offers a promising solution for achieving data privacy by enabling secure computations directly on encrypted data. In this paper, aimed at computational scientists, we explore the viability of FHE-based, privacy-preserving numerical simulations of partial differential equations. We begin with an overview of the CKKS scheme, a widely used FHE method for computations with real numbers. Next, we introduce our Julia-based packages OpenFHE.jl and SecureArithmetic.jl, which wrap the OpenFHE C++ library and provide a convenient interface for secure arithmetic operations. We then evaluate the accuracy and performance of key FHE operations in OpenFHE as a baseline for more complex numerical algorithms. Following that, we demonstrate the application of FHE to scientific computing by implementing two finite difference schemes for the linear advection equation. Finally, we discuss potential challenges and solutions for extending secure numerical simulations to other models and methods. Our results show that cryptographically secure numerical simulations are possible, but that careful consideration must be given to the computational overhead and the numerical errors introduced by using FHE.
@@ 1 Introduction
Partial differential equations (PDEs) are widely used to model phenomena across nearly every scientific field, from physics and engineering to biology and finance. Since many PDEs cannot be solved analytically, numerical methods are commonly used to approximate their solutions. Beyond strictly scientific settings, PDE-based simulations also play a key role in industry and business applications. In certain cases, the input data and simulation results involve highly sensitive information—–such as personal health records, financial details, or proprietary engineering designs–—that must be carefully protected. For instance, finite element simulations of the aorta used in personalized approaches for treating cardiovascular disease rely on patient-specific data [1, 2]. These datasets must be handled securely to comply with regulations like the General Data Protection Regulation (GDPR) in the European Union or the Health Insurance Portability and Accountability Act (HIPAA) in the United States, posing challenges for scientific research as well as practical implementations [3, 4]. Another example arises in population genetics, where the evolution of gene frequencies has long been modeled by reaction-diffusion equations [5, 6, 7]. Due to nonlinear dynamics and complex boundaries, these models are typically not amenable to analytical solutions and are therefore solved numerically [8, 9, 10]. A practical issue here is working with combined datasets from multiple providers (which do not trust each other), while still protecting the privacy of individual genetic information [11, 12, 13].
To safeguard data, encryption is commonly employed both at rest and during transmission. However, once the data is decrypted for processing, it becomes vulnerable to attacks. The data privacy risks associated with decryption for processing become especially pronounced when data must be processed remotely, as is often the case with public cloud computing platforms. Limited local resources or the need for distributed systems may necessitate offloading computations to external cloud infrastructures, thereby exposing decrypted data to potential attacks. With the rapid expansion of cloud computing [14], ensuring data privacy during computation has emerged as a pressing concern for both practitioners and researchers [15, 16].
@@
In this paper, we aim to address these issues by examining fully homomorphic encryption from a scientific computing perspective and introducing it to the computational physics community. We begin with a brief overview of the CKKS scheme for computational scientists, and present two FHE packages we developed in the Julia programming language that allow one to easily experiment with novel FHE applications. We then use these packages to analyze the performance and accuracy of basic FHE operations to get a baseline for more complex algorithms. Finally, we will demonstrate the first prototypes of FHE-secured numerical simulations by implementing and evaluating two well-known numerical methods for solving partial differential equations. We will find that while secure numerical simulations through FHE are indeed possible, the computational overhead is currently significant for many practical scenarios, and the numerical errors introduced by the encryption techniques must be carefully considered when designing applications. All code used in this paper, as well as the data generated during the experiments, is available in our reproducibility repository [30].
The remainder of the paper is structured as follows: In Sec. 2, we give an overview of the CKKS scheme for FHE with real numbers. Moreover, we introduce the open-source C++ library OpenFHE and the Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a convenient interface for secure arithmetic operations in Julia for experimentation and rapid prototyping. We examine the implementation of the CKKS scheme in OpenFHE in Sec. 3, and analyze the performance and accuracy of basic FHE operations. In Sec. 4, we implement several finite difference schemes to solve the linear advection equation using FHE, and in Sec. 5 we analyze the performance and accuracy of the cryptographically secured numerical solvers. Sec. 6 explores challenges and potential solutions for extending secure numerical simulations to other mathematical models and numerical methods. Finally, we summarize our findings in Sec. 7 and provide an outlook on future work.
@@ 5.4 Parallel computations
6 Extension to other mathematical models and numerical methods
As demonstrated in the previous section, secure numerical simulations are generally possible. It was also noted that the primary impediments to the practical use of the CKKS scheme for such scenarios are the large computational overhead and, to a lesser extent, the precision loss. Apart from this, three main challenges arise when considering to extend the simulation prototypes shown in this paper to other PDEs and to more complex numerical schemes: a single built-in data structure, limited function evaluation capabilities, and constraints in control flow logic when it comes to conditional branching. In the following, we will discuss these issues and potential solutions to them, to give other researchers a starting point for their own investigations into secure numerical simulations. We will conclude this section with a brief summary of some general strategies for improving performance.
6.1 Data structures
A vector of real numbers is the only encrypted data structure available in the CKKS scheme, as described in Sec. 2.1. More complex data structures, such as tensors or matrices, are not natively supported. In Sec. 4.3.2, we have introduced one potential approach to a matrix representation for use with explicit, mesh-based numerical simulations. In other algorithms that involve matrix-matrix or matrix-vector multiplications, specialized encoding strategies for the matrix lead to more efficient algorithms in terms of the multiplicative depth [55, 56, 28].
In general, the concrete choice of data representation depends on the target application and the used numerical method. Due to the significant computational overhead of the CKKS scheme, it is necessary to make full use of its batch processing capabilities, i.e., the ability to do element-wise addition and multiplication on the entire data vector. Therefore, discretization schemes that allow for efficient vectorization are preferable when considering other numerical methods for secure simulations. The choice of PDE is largely unaffected by the data structure, as long as a single model with a constant number of state variables is used throughout the computational domain.
6.2 Function evaluation
As discussed before, the CKKS scheme supports only a limited set of arithmetic operations, i.e., element-wise addition and multiplication of the ciphertexts, as well as rotation. This means that more complex arithmetic operations, such as division, exponentiation, or trigonometric functions, are not directly available but must be approximated by a series of simpler operations or using a polynomial interpolation. For example, the division of two ciphertexts can be implemented with the Newton-Raphson method [35] or Goldschmidt’s algorithm [36], the latter of which is also used to implement floating point division in some CPU families [57]. Exponentiation or trigonometric functions may be approximated by a Taylor series expansion [58].
Since evaluating a polynomial expression is straightforward with the built-in arithmetic operations, any function that is reasonably well approximated by a polynomial can be computed with sufficient accuracy. For many other functions, an iterative approach is required. Here, the main difficulty is that the number of iterations must be fixed a priori, since branching based on encrypted data is not possible (refer to the next section for details). Without fundamental limitations to computing branch-free functions with FHE, the primary challenge is therefore to find an efficient approximation that converges robustly. Selecting the best approach is problem-dependent and likely requires some experimentation to find a good balance between accuracy and performance. When considering other mathematical models or numerical discretizations, it is thus advisable to choose options that minimize the multiplicative depth of the FHE operations to avoid costly bootstrapping operations.
For many discontinuous functions, e.g., sign/comparison, the use of CKKS is challenging as the polynomial approximation in the proximity of the singularity point(s) requires extremely high-degree polynomials, i.e., very large multiplicative depth and runtime. Instead, other FHE schemes are often used, such as DM [23] or CGGI [24]. DM and CGGI support the evaluation of arbitrary functions via lookup tables using a technique called functional or programmable bootstrapping [59]. In these scenarios, a scheme switching procedure is employed [60, 61], where the encryption of the data is temporarily changed from CKKS to DM/CGGI. This allows for the evaluation of the discontinuous function, after which the data is converted back to a CKKS representation. Note that such scheme switching capability is supported in OpenFHE [40]. The main drawback of scheme switching is the significant computational overhead, especially when high precision is required [62, 51, 63]. A promising new method for evaluating discontinuous functions using functional CKKS bootstrapping is proposed and implemented in [64]. The method enables arbitrary function evaluation via lookup tables and achieves a throughput that is orders of magnitude higher than DM or CGGI.
6.3 Conditional branching
By their very nature, many of the control flow directives commonly used in regular programs cannot be employed with FHE methods. More specifically, FHE precludes the use of conditional statements or branches—including loops with a termination condition—that depend on encrypted data. This is due to the fact that the encryption of the data prevents the processor from knowing the actual values of the data. If this were not the case, it would be possible for a malicious actor to infer information about the encrypted data, e.g., by using a bisection algorithm to compare the secret ciphertext with known values.
As noted in [65], this restriction in control flow logic presents a fundamental challenge when converting regular programs to FHE algorithms. Since branching on an encrypted value is impossible by design, an FHE algorithm encountering a conditional branch must therefore first compute the outcome for each possible branch. The final result is then constructed by combining the results from all branches and applying an appropriate mask that selects the correct result. For example, Alg. 4 shows a simple branching statement in a regular program. Based on a function condition() that returns either $1$ or 00, the value of $r$ is computed by calling either foo() or bar(). The equivalent FHE algorithm in Alg. 5 first computes the value of the condition, evaluates both branches, and then combines the results with the condition value to obtain the final result.
Algorithm 4 Regular conditional branching.
[…]
1:$r\leftarrow\text{foo}()$Else
2:$r\leftarrow\text{bar}()$EndIf
[…]
Algorithm 5 Conditional branching with FHE.
1:$c\leftarrow\text{condition()}$()
2:$r_{\text{foo}}\leftarrow\text{foo}()$
3:$r_{\text{bar}}\leftarrow\text{bar}()$
4:$r\leftarrow c*r_{\text{foo}}+(1-c)*r_{\text{bar}}$
Such an approach is feasible if a small, predetermined number of branches needs to be considered, e.g., when computing the well-known HLLC flux for the compressible Euler equations, which has different branches based on three different wave speeds [66]. It becomes intractable, however, if the number of branches is large and possibly unknown in advance, e.g., when terminating a time step loop while using a dynamic, CFL-based time step size. Furthermore, numerical schemes that strictly require conditional branching based on current solution values are not directly compatible with FHE. This includes methods that dynamically modify their data structures during the simulation, such as periodic remeshing when dealing with deforming boundaries or adaptive mesh refinement. As a result, this poses a concrete limitation on the applicability of FHE to certain classes of numerical simulations.
Therefore, when choosing a numerical method for secure numerical simulations, it is important to avoid the need for control flow that depends on encrypted data and that does not have a fixed number of branches. While the statement that FHE can be used to compute any function (see Sec. 2) is not violated, it means that in practice, the translation to an FHE algorithm is not always feasible. As a workaround, one can consider allowing certain control flow operations to be unencrypted, like we did in Sec. 4.4 for the time loop in Alg. 3.
6.4 General strategies for improving performance
As noted in the previous subsections, most methods for dealing with mathematical models and numerical schemes that require more complex FHE operations incur a significant computational overhead. To mitigate this, several strategies can be employed to optimize the performance of FHE algorithms:
  • 1.
    Prefer parallelizable algorithms over sequential ones. As discussed in Sec. 2.1.6, using a binary approach for multiple multiplications can reduce the overall multiplicative depth of the FHE operations. This naturally applies also to all other functions that consume multiplicative levels. For example, consider loop unrolling to manually encode its iterations more efficiently.
  • 2.
    Minimize or avoid FHE-hard operations. For instance, another set of state variables, e.g., using a different non-dimensionalization strategy, may save some (costly) division operation in each time step.
  • 3.
    Reduce the number of rotations. While they do not consume multiplicative levels, they require similar computational resources as multiplications (see Table 2). For example, summation over a row/column of a matrix may be expressed as a series of additions instead.
  • 4.
    Consider performing costly computations a priori. For example, it might be more efficient to use a fixed, smaller time step based on an conservative estimate, rather than computing a potentially larger step size dynamically in each time step loop iteration.
A more detailed description of these and other optimization strategies can be found in [28].
@@ 7 Conclusions and outlook
Next, we implemented the first-order upwind and second-order Lax-Wendroff schemes for the linear advection equations for secure numerical simulations using the SecureArithmetic.jl package. For this purpose, we introduced a $\operatorname{circshift}$ operation that is compatible with the CKKS scheme and that works for both 1D and 2D data representations. Convergence tests verified the correctness of our implementations and showed that the CKKS encryption does not break the convergence properties of the finite difference schemes.
@@
We then conducted homomorphically encrypted numerical simulations by building upon the tools and methods developed in the previous sections. Our results illustrate that it is feasible to implement secure numerical simulations with the CKKS scheme, with results that match the unencrypted simulations well. However, the encrypted operations introduce a large computational overhead and are a significant source of error, to the point that we were able to achieve less than single-precision floating point accuracy at the cost of a six-orders-of-magnitude increase in runtime. Nearly all of these undesirable effects can be attributed to the bootstrapping procedure, which is the most inaccurate and slowest operation in the CKKS scheme. However, bootstrapping is also necessary to refresh the multiplicative depth of the ciphertexts, making it indispensable for long-running simulations. Moreover, iterative bootstrapping can further improve the accuracy of an FHE algorithm, though at a non-negligible runtime penalty that needs to be taken into account. We also showed that the CKKS scheme implementation in OpenFHE is parallelizable, and that using multiple threads can lead to a significant reduction in the solution time.
Finally, we explored challenges—and how to address them—when extending secure numerical simulations beyond the presented examples. Many of the discussed issues can be mitigated through careful algorithm design and by accepting some computational overhead. However, the lack of dynamic, value-based conditional branching precludes the use of schemes that strictly depend on it, such as adaptive mesh refinement.
@@
To summarize, our work demonstrates that it is possible to perform secure numerical simulations using fully homomorphic encryption. At the same time, the large computational overhead and precision loss in the bootstrapping operation are significant obstacles that need to be taken into account. Therefore, currently FHE-secured numerical simulations only make practical sense in specialized applications where the security of the data is of utmost importance. Thus, more detailed investigations are necessary to optimize the use of the CKKS scheme for numerical simulations. By providing a user-friendly interface to the OpenFHE library in Julia, and by implementing some important building blocks such as the $\operatorname{circshift}$ algorithms, we hope to lay the foundation for further research in this direction. In addition, our discussion of the challenges and potential solutions for extending the secure numerical simulations to other mathematical models and numerical methods should provide a starting point for other researchers interested in this field.
@@
In the future, we plan to extend our work in several areas. We would like to increase the complexity of the simulations by looking at systems of equations and at three-dimensional problems. This will also necessitate an extension of the $\operatorname{circshift}$ operation to support multiple ciphertexts and tensors. Furthermore, we will investigate the discretization of nonlinear equations, which will require us to handle additional arithmetic operations such as division and exponentiation. Moreover, we would like to better understand the observed behavior that the accuracy of secure numerical simulations with bootstrapping can improve over time. Finally, we plan to continue developing the SecureArithmetic.jl package as a tool for experimentation and rapid prototyping with encrypted algorithms, to open up FHE to a wider audience in the computational science community.
@@ Acknowledgments
We gratefully acknowledge the computing time provided by Numerical Simulation research group of Gregor Gassner, University of Cologne. Michael Schlottke-Lakemper further acknowledges funding through the DFG research unit FOR 5409 "Structure-Preserving Numerical Methods for Bulk- and Interface Coupling of Heterogeneous Models (SNuBIC)" (project number 463312734), as well as through a DFG individual grant (project number 528753982). Moreover, we would like to thank Andrew Hazel for his valuable feedback on the manuscript.